g={(-6,-5),(2,-6),(8,-2),(9,8)}
h(x)=2x+13
Find the following
g^-1(-6)=
h^-1(x)=
(h°h^-1)(9)=

Answer by CubeyThePenguin(3113)

g^(-1)(-6) = what point has y-coordinate (-6) = 2

y = 2x + 13
(y-13)/2 = x

h^(-1)(x) = (y-13)/2

h(h^(-1)(9)) = 9